OpenStudy (slayer8001):

3(5 − 2x) = −2(6 – 3x) − 10x 1. x=38 2.x=-37/38 3.x=-260 4.x=13.5

OpenStudy (anonymous):

g Let's solve your equation step-by-step. 3(5−2x)=−2(6−3x)−10x Step 1: Simplify both sides of the equation. 3(5−2x)=−2(6−3x)−10x Simplify: (Show steps) −6x+15=−4x−12 Step 2: Add 4x to both sides. −6x+15+4x=−4x−12+4x −2x+15=−12 Step 3: Subtract 15 from both sides. −2x+15−15=−12−15 −2x=−27 Step 4: Divide both sides by -2. −2x −2 = −27 −2 x= 27 2 Answer: x= 27 2